OpenStudy (anonymous):

what is the arcsin(sin3pi)

zepdrix (zepdrix):

The range of arcsine is restricted from \(\large -\pi/2\) to \(\large \pi/2\). \[\large \arcsin\left(\sin(3\pi)\right)\qquad = \qquad \arcsin\left(0\right) \qquad = \qquad 0\]

OpenStudy (anonymous):

Those domain restrictions always find a way to ruin your day if you neglect them.

OpenStudy (anonymous):

darn. and it promptly did